Southwestern Fashions, Inc. which uses a job-order costing system, had two jobs in process at the start of the year: job no. 101 ($84,000) and job no. 102 ($53,300). The following information is available:

  
a.

The company applies manufacturing overhead on the basis of machine hours (based on practical capacity). Budgeted overhead and machine activity for the year were anticipated to be $824,000, and 16,000 hours, respectively.

b.

The company worked on four jobs during the first quarter. Direct materials used, direct labor incurred, and machine hours consumed were as follows:

Job No. Direct Material Direct Labor Machine Hours
101 $21,000 $35,000 1,200   
102 22,000 700   
103 44,000 65,000 2,000   
104 15,000 8,800 500   
c.

Manufacturing overhead during the first quarter included charges for depreciation ($33,000), indirect labor ($60,000), indirect materials used ($5,200), and other factory costs ($139,600).

d.

Southwestern Fashions completed job no. 101 and job no. 102. Job no. 102 was sold on account, producing a profit of $34,800 for the firm.

4. Did the finished-goods inventory increase or decrease during the first quarter? By how much?
5. Was manufacturing overhead under- or overapplied for the first quarter of the year? By how much?
Answer #1

Manufacturing overhead rate = 824,000/16,000 = $51.5 per machine hour

Job No. Balance at Beg. of year (A) Direct Material (B) Direct Labor (C ) Manufacturing Overheads
Machine Hours X $51.5 (D)
Cost of Finished Goods (A+B+C+D) = E Cost of Goods Sold (F) Closing Finished Inventory (E-F) = G
101                84,000                      21,000                                35,000                        61,800                             201,800                                       -                      201,800
102                53,300                               -                                  22,000                        36,050                             111,350                            111,350                               -  
103                         -                        44,000                                65,000                      103,000 Not Completed Not Completed Not Completed
104                         -                        15,000                                  8,800                        25,750 Not Completed Not Completed Not Completed
Total             137,300                      80,000                             130,800                      226,600                             313,150                           111,350                    201,800

4. Finished inventory has Increased during first quarter by $64,500 , ie 210,800 (G) - 137,300 (A)

5. Budgeted overheads applied in first quarter = $226,600 (D)

Actual overheads for the quarter = 33,000 + 60,000 + 5,200 + 139,600 = $237,800

Actual overheads rare higher than budgeted overheads by $11,200 resulting in under application.